solve e^x/9-8=6. Add 2.248 to the value of x. Round to the nearest whole number.
Answer by stanbon(75887)

solve e^x/9-8=6.
----
Solve for "x":
e^(x/9) = 14
----
Take the natural log to get:
x/9 = ln(14)
----
x = 9*ln(14)
x = 23.7515
===================
Add 2.248 to the value of x. Round to the nearest whole number.
x = 23.7515 + 2.248
x = 25.99
Rounded: x = 26
===================================
Cheers,
Stan H.
===================================